# A mechanism-informed reinforcement learning framework for shape optimization of airfoils

### v1

@@ 4.3 Mesh deformation for unstructured mesh
Figure 8: Mesh around the tail before curvature capture; Mesh around the tail after curvature capture.
To calculate the curvature for each boundary curve, we shall find the parameters correspond with the endpoints, $t_{1}=B^{-1}(\mathbf{x}_{1})$ and $t_{2}=B^{-1}(\mathbf{x}_{2})$. Then we generate sample points that are dense enough within such segment, $sp_{0},sp_{1},\dots,sp_{L+1}$. For three consecutive sample points $sp_{i-1},sp_{i},sp_{i+1}$, we use $\theta_{i}$ to denote the angles between $\overrightarrow{sp_{i-1}sp_{i}}$ and $\overrightarrow{sp_{i}sp_{i+1}}$, then $\kappa_{i}={\theta_{i}}/{\|\overrightarrow{sp_{i}sp_{i+1}}\|}$ is used to denote the discrete curvature values. $\sum_{i=0}^{L}\kappa_{i}$ is used to denote the curvature change within the segment. It is shown in the left part of Figure 8, that if the curvature change scale is smaller than the mesh size like the elements near the tail, the mesh can not detect such variation. However, when we use elements with smaller sizes, it can capture the curvature change within the segment. After running this algorithm for several times, the mesh can depict a more elaborate airfoil shape.
4.4 Curve smoothing
The reinforcement learning process in this framework necessitates a delicate balance between the granularity of the sample points and the computational feasibility. High-dimensional state representations, while offering detailed control over the geometry, exacerbate the curse of dimensionality. This not only makes the optimization landscape more complex and harder to navigate but also strains the underlying PDE solvers. For machine learning-based solvers, deviations from the trained geometries can result in diminished accuracy, adversely affecting the fidelity of the objective functional evaluations. This necessitates the integration of traditional numerical methods capable of handling a wider variety of shapes with robustness, ensuring high precision in objective functional computation irrespective of the geometric configuration.
Furthermore, excessive sample points can introduce local singularities and discontinuities in the geometry, making it highly non-smooth such as shown in the left part of Figure 9. While actual engineering solutions rarely embody such erratic geometric characteristics, their presence during the learning process can significantly affect the optimization. Non-smooth geometries not only challenge the solver’s capacity to provide accurate solutions but also obscure the learning signal, making it difficult for the agent to discern productive learning directions. Such geometric irregularities could mislead the agent, leading to suboptimal policies that fail to converge to the desired outcome.
@@ 4.4 Curve smoothing
Figure 9: Left: mesh around the shape without curve smoothing; Right: mesh around the shape with the curve smoothing;
A refined approach entails leveraging dimensionality reduction and feature engineering to extract and condense essential geometric features into a more succinct and computationally efficient representation. While directly reducing the dimensionality of sample points may compromise their utility for precise geometric manipulation, alternative strategies can offer a balance between detail and manageability. The utilization of the Bézier curve fitting adopted above stands out as a proficient technique for reducing dimensionality while capturing critical curve characteristics. This method effectively summarizes vital information, facilitating a streamlined representation of curves. However, the challenges still persist, particularly regarding the smoothness of the curves. Merely lowering the order of Bézier curves can inadvertently eliminate valuable local details. To circumvent these issues, incorporating smoothing techniques during the Bézier curve generation process emerges as a potent solution. Integrating regularization directly into the curve fitting algorithm ensures the smoothness of the curves, preserving the integrity of local geometric features without sacrificing the overall quality of the representation.
The Bézier curve fitting adopted above is :
$\begin{array}[]{cc}&\text{Find}\qquad\{CP_{i}\}_{i=0}^{n},\\ &\min\qquad\sum_{j=0}^{L+1}|\!|B(t_{j})-sp_{j}|\!|^{2}\end{array},$ (23)
where $t_{j}$ is the parameters corresponding to the sample points $sp_{j}$. With the regularization term added to the system, the equation becomes:
@@
where $\lambda_{s}$ stands for the smoothing coefficient. The impact of this regularization term is to make the control points distributed equidistantly. As a result, the potential issues of the non-smooth curve can be mitigated effectively. As shown in the right part of Figure 9, the deformed curve preserves a satisfactory geometry while having a detailed control of the airfoil.
5 Design of Twin Delayed DDPG
In the previous section, we have established a simple framework that can manipulate the deformation of the airfoil shape. However, training an agent to achieve the goal of interacting with the environment in such a complicated process is not straightforward. Firstly, the dimension of the state is defined by hundreds of points. During the initial stage, the difference between the tensor representation of different airfoils is tiny. Then the agent may not detect the features of different airfoils. Secondly, the inherent relation between the state and action with corresponding action parameters is intricate. Since the authentic value of reward needs the calculation with a CFD solver, the critic neural network is adopted to approximate such a process. The neural network of reinforcement learning is usually constructed with a light structure since the sampling is expensive, and the result shall be obtained quickly for a long-term iteration. Conversely, a simple neural network is not accurate enough to capture the relation between these variables. Meanwhile, a complicated neural network suffers a lot from the gradient disappear problem, leading to the training invalid. Thirdly, during the training process, the agent may fail to learn the optimization method if it can not obtain a positive reward. How to guide this agent to take appropriate actions on the airfoil is crucial for the policy update. Last but not least, the DDPG framework suffers a lot from the overestimate. With the development of an algorithm in reinforcement learning, the Twin Delayed DDPG framework remits such influence efficiently and enhances the training performance. In this section, we focus on these issues and elaborate on the details of the Twin Delayed DDPG framework for the design of airfoil.
5.1 Actor and critic
@@
The actor is developed to choose the potential action based on the input state such that the obtained reward is maximal. While the state is the tensor representation of the corresponding airfoil shape, the actor should capture the main feature and distinguish different shapes. Aiming at having a precise control of the airfoil, the geometry is constructed with hundreds of points pair. However, the agent may fail to learn valuable actions based on different states when this tensor representation is conducted within a high-dimensional space. A new state conducted by a specific action only takes a small change of target location. Such difference is easy to overlook within a hundred-dimensional tensor. The smooth action process explained in the previous section can remit such influence but the essential difficulties have not been solved. Trying to conduct the deformation within a lower dimensional tensor space, we considered adopting the variational autoencoder(VAE) module at first.
  • Variational autoencoder
    The module contained a neural network that is designed to capture complex relationships between different airfoils, which is crucial for accurately modeling the state space in the wing design problem. The VAE will help in generating new, valid states that conform to the physical constraints of the wing design, aiding in exploring the state space more effectively. The encoder part of VAE maps the variable to a distribution in the latent space, rather than a single point. This is typically a Gaussian distribution characterized by vectors of means $\mu_{vae}$ and log-variances $\sigma_{vae}$. It is capable of learning the generative process for complex data. In this context, by learning how existing wing shapes are generated, a VAE can help in determining the actions to take for optimizing new designs.
    To guarantee the latent layer captured the key feature of airfoil data $\mathcal{X}$, the decoder part will reconstruct the latent representation to $\widetilde{\mathcal{X}}$. The standard loss function for the VAE is a combination of two distinct terms: the reconstruction loss and the Kullback-Leibler (KL) divergence[58]. The reconstruction loss measures the difference between the original input data $\mathcal{X}$ and the reconstructed data $\widetilde{\mathcal{X}}$ produced by the decoder of the VAE. It ensures that the VAE is capable of accurately reconstructing inputs, which means that the latent space is capturing the necessary information about the data distribution.
Layer Type Layer Name Operation Output Channels Activation
Input input - 15 * 2 * 2 -
Actor
nn.LayerNorm norm1 Layer Normalization 256 -
ResidualBlock res_block1 Residual Block 256 -
nn.Linear fc Linear 512 SELU
nn.LayerNorm norm2 Layer Normalization 512 -
nn.Linear layer_action_param Linear 128 SELU
Postprocessing
nn.Linear action_probability Linear 3 softmax
nn.Linear action_param Linear 5 softsign
Table 1: Actor Network Architecture
The variational autoencoder meets the requirement of dimensionality reduction efficiently. However, it brings extra problems to the construction of the learning process. Firstly, it is not easy to take the balance of the reconstruction error and original loss function. Secondly, deeper neural networks aggravate the risk of the gradient disappearance problem. Thirdly, the sampling process is expensive in this reinforcement learning task. The VAE module can not perform well with limited data and gets trapped into overfitting issues. Additionally, failures of this model may be attributed to the reasons as follows. If we collect the $266$-dimensional tensors representing the airfoil, each tensor can be treated as a point within this high-dimensional space. Specifically, points that conform to closed-curve airfoil geometries form a submanifold within this space. This submanifold is non-convex since states between distinct airfoil shapes may include mesh tangling form. The probability that these points are collected is 00 since we avoid such circumstances with the methods mentioned above. The challenge of dimensionality reduction in this context is underscored by the difficulty in preserving geometric and topological properties during the mapping to a lower-dimensional space, a process complicated by the non-convex nature of the submanifold. As a result, the VAE model faces significant hurdles in accurately representing such complex, non-convex geometries without incurring substantial reconstruction errors. The optimal transportation method proposed in [59] aims at finding a measure-preserving map that minimizes the total transport cost. However, it may not be practical for our current framework to conduct analysis with the collected data in a large magnitude.
@@
Then we tried a different method, using Bézier curves representation for the dimensionality reduction. The Bézier curves take the most important information for specific curves. Besides, the difference between different control points originating from corresponding states is amplified with the least square process. It helps the agent to distinguish different states. Meanwhile, as an independent process, it avoids the gradient calculation within a neural network.
Layer Type Layer Name Operation Output Channels Activation
Input input - 66 * 2 * 4 -
Attention Module
nn.linear state2hidden Linear 64 SELU
nn.linear action2hidden Linear 64 SELU
nn.linear param2hidden Linear 64 SELU
nn.MultiheadAttention self_attn Self-Attention 64 -
nn.MultiheadAttention cross_attn Cross-Attention 64 -
Critic processing
ResidualBlock res_blockCritic1 Residual Block 512 SELU
ResidualBlock res_blockCritic2 Residual Block 128 SELU
nn.Linear fusion_fc2 Linear 8 SELU
nn.Linear q_value_fc Linear 1 -
Table 2: Critic Network Architecture
@@
The dimension is reduced efficiently with the Bézier curve representation. Then the information should be processed through the neural network constructed as Table 1 to obtain the action that maximizes the reward from the current state. It should be noted that the output includes action probability for different types of action and corresponding parameters that action possesses. The action that is chosen to be performed originates from the $softmax$ layer while the corresponding parameters are based on the restriction of shape design. For fear of mesh tangling, an activation function needs to be designed that maps the parameters to reasonable intervals. However, the traditional activation functions like $tanh$ and $sigmoid$ suffer a lot from the gradient disappear phenomenon. Then the $softsign$ type activation function is adopted. Furthermore, the residual connection module is also integrated to guarantee the gradient can be propagated normally within the neural network.
@@
The critic neural network is designed to evaluate the potential feedback for various actions acted on different states based on the experience accumulated in the replay buffer. The reward mechanism is complicated since the data originate from a CFD solver, not even to mention the shape deformation. In order to approximate such a process, the interaction between the shape, actions and corresponding parameters should be modeled efficiently. We considered adopting the self-attention and cross-attention mechanism together to investigate the inherent relation. The self-attention mechanism can detect those features which take a significant influence on the target functional. Meanwhile, the cross-attention mechanism is adopted to approximate the shape deformation process. In other words, it tries to understand how the action with corresponding parameters influenced the shape deformation. The state, action, and action parameters are mapped to the hidden layer with the same dimension to ensure the cross-attention layer can be modeled.
@@
The loss function is an important component of the learning framework which decides the update direction of parameters in neural networks. For different modules in reinforcement learning, the loss function is defined according to the features. In order to choose the action that maximizes the reward based on specific states, the agent should evaluate the potential feedback correctly at first. It indicates that the critical neural network should be updated more frequently than the actor neural network. The long-term reward loss is defined as
@@ 5.2 Exploration and exploitation
Exploration and exploitation are fundamental components of the interaction process. The agent tries different actions to gain information about the environment during the exploration process. Extensive exploration is crucial since the agent needs to try different actions to gain more information about the environment within an unfamiliar task. It helps avoid local optima and enables the agent to find more effective strategies. Then the agent uses its existing knowledge or experience to make decisions that maximize immediate rewards during the exploitation period. Balancing exploration and exploitation is vital in reinforcement learning. Excessive exploration can prevent the agent from fully utilizing known effective strategies. On the other hand, too much exploitation might cause missing out on better strategic options. In this framework, we adopt the $\epsilon$-greedy strategy. This is a popular technique that chooses an action at random with probability $\epsilon$ and selects the action that is predicted to have the highest reward based on its current knowledge with probability $1-\epsilon$. Besides, in our framework, we integrate Ornstein-Uhlenbeck (OU) noise as a complementary tool to the $\epsilon$-greedy strategy, specifically targeting the exploration aspect of reinforcement learning in continuous action spaces. While the $\epsilon$-greedy strategy efficiently balances exploration and exploitation by randomly selecting actions with a probability $\epsilon$, OU noise adds an additional layer of sophistication to the exploration process.
@@
Even though the optimization mechanism can be well implemented with the exploration and exploitation process, the learning effect is not ideal usually. The difficulties originate from the complex mechanisms involved in airfoil design. Firstly, it is not easy to identify the correct reward direction in random exploration. Excessive penalization can lead the neural network parameters into an unfavorable bottleneck, where the agent opts for minimal action to avoid penalties, eventually leading to a suboptimal local optimum. Secondly, airfoil design inherently demands a geometry with a certain level of smoothness. If the shape changes drastically in a single step, even a correct design trend may result in unfavorable reward feedback due to geometric roughness. This also indicates that the shape optimal design needs to be implemented within an elaborate parameter representation environment. Thirdly, the training batch size is small during the initial stage since the available data is limited. A smaller batch size can lead to more noise in the gradient estimates, which can inadvertently aid exploration by preventing the model from quickly converging to a local optimum. However, this noise can also result in less stable learning, as the updates might oscillate or diverge.
@@
To address the above issues, we introduced several improvement strategies as follows, i). if a certain action results in a penalty, the agent attempts to explore the opposite direction starting from the pre-deformation geometry. Both exploration results and parameters are stored in the replay buffer, and the action leading to a better outcome is chosen for the next step. This mechanism significantly reduces aimless, detrimental exploration in airfoil design, preventing the model from extensively exploring in unproductive directions. This is also inspired by the idea of a twin network to prevent oscillation during the simulation and it turned out to be very effective for reinforcement learning. Similarly, if the current action brings a perfect reward, the agent will be encouraged to take a similar action with a tiny noise. As the environment is not straightforward to obtain positive meaningful rewards, such a mechanism can help the agent to accumulate useful experience quickly. ii). the single step movement of the sample points is constrained, so that there may not occur singular geometry. Besides, an early stop mechanism is constructed when the accumulated reward decreases to the minimum threshold. This can be explained as the agent failing for the current design process when the geometry gets trapped into a very terrible shape and may not recover easily. Then the environment should be reset for the next turn. iii). the training process is separated into different periods. A small batch size is adopted for the initial stage. When the replay buffer accumulates enough data, a larger batch size number will be updated. Besides, the data are stored in the replay buffer with a 4-tuple form transition $(state,action,action\_parameters,next\_state)$. Every time the gradient should be calculated from the 4-tuple with a number of batch sizes. The components include three parts, the recent transitions, the transitions with the best reward and randomly chosen transitions from the whole memory. The recent transitions represent the behavior under the current parameters of neural networks while the best transitions can help the agent to repeat the beneficial actions and the random transitions can make the gradient calculation more stable. With such efficient exploration and exploitation mechanisms, we step further to fulfill the reinforcement learning framework.

### v2

@@ 3.3 Action of mesh deformation on unstructured mesh
Figure 5: Mesh around the tail. Left: before curvature capture; Right: after curvature capture.
To calculate the curvature for each boundary curve, we shall find the parameters correspond with the endpoints, $t_{1}=B^{-1}(\mathbf{x}_{1})$ and $t_{2}=B^{-1}(\mathbf{x}_{2})$. Then we generate sample points that are dense enough within such segment, $sp_{0},sp_{1},\dots,sp_{L+1}$. For three consecutive sample points $sp_{i-1},sp_{i},sp_{i+1}$, we use $\theta_{i}$ to denote the angles between $\overrightarrow{sp_{i-1}sp_{i}}$ and $\overrightarrow{sp_{i}sp_{i+1}}$, then $\kappa_{i}={\theta_{i}}/{\|\overrightarrow{sp_{i}sp_{i+1}}\|}$ is used to denote the discrete curvature values. $\sum_{i=0}^{L}\kappa_{i}$ is used to denote the curvature change within the segment. It is shown in the left part of Figure 5, that if the curvature change scale is smaller than the mesh size like the elements near the tail, the mesh can not detect such variation. However, when we use elements with smaller sizes, it can capture the curvature change within the segment. After running this algorithm several times, the mesh can depict a more elaborate airfoil shape. Even though issues remains on the irregular shapes as we try to manipulate the airfoil with hundreds of sample points.
• Irregular shape. High-dimensional state representations, while offering detailed control over the geometry, exacerbate the curse of dimensionality. This not only makes the optimization landscape more complex and harder to navigate but also strains the underlying PDE solvers. For machine learning-based solvers, deviations from the trained geometries can result in diminished accuracy, adversely affecting the fidelity of the objective functional evaluations. This necessitates the integration of traditional numerical methods capable of handling a wider variety of shapes with robustness, ensuring high precision in objective functional computation even when the configuration changed significantly.
Furthermore, hundreds of sample points can introduce local irregularity to the geometry such as shown in Figure 6. While actual engineering solutions rarely embody such erratic geometric characteristics, their presence during the learning process can significantly affect the optimization. Irregular geometries not only challenge the solver’s capacity to provide accurate solutions but also obscure the learning signal, making it difficult for the agent to discern productive learning directions. Such geometric irregularities could mislead the agent, leading to suboptimal policies that fail to converge to the desired outcome.
@@
Figure 6: Meshes around the airfoil. Left: without curve smoothing; Right: with curve smoothing;
An improved approach entails adopting dimensionality reduction to extract essential geometric features into a more computationally efficient representation. While directly reducing the dimensionality of sample points will compromise their utility for precise geometric manipulation, strategies should be designed to take a balance between the local controllability and regularity. The utilization of the Bézier curve fitting stands out as a proficient technique for reducing dimensionality while capturing critical curve characteristics. This method effectively summarizes vital information, facilitating a streamlined representation of curves. However, the challenges still persist, particularly regarding the regularity of the curves. Directly lowering the order of Bézier curves can inadvertently eliminate valuable local details. To circumvent these issues, incorporating smoothing techniques during the Bézier curve generation process emerges as a potent solution. Integrating regularization directly into the curve fitting algorithm ensures the regularity of the curves, preserving the integrity of local geometric features without sacrificing the overall quality of the representation.
The Bézier curve fitting for above purpose can be formulated by :
$\begin{array}[]{cc}\text{Find}\quad\{CP_{i}\}_{i=0}^{n},\qquad\min\quad\sum_{j=0}^{L+1}|\!|B(t_{j})-sp_{j}|\!|^{2},\end{array}$ (16)
where $t_{j}$ is the parameters corresponding to the sample points $sp_{j}$. With a regularization term added to the system, the equation becomes:
@@
where $\lambda_{s}$ stands for the smoothing coefficient. Such an regularization term is to make the control points equally distributed. Then the potential issues of the irregular curve can be mitigated effectively. As shown in 6, the deformed curve preserves a satisfactory geometry while having a detailed control of the airfoil.
4 Design of Twin Delayed DDPG algorithm
In the previous section, we introduced the learning process that can manipulate the deformation of the airfoil shape. However, training an agent to achieve the goal of interacting with the environment in such a complicated process is not straightforward. With the development of algorithm in reinforcement learning, the Twin Delayed DDPG framework enhances the training performance effectively. In this section, we will elaborate on the details of the Twin Delayed DDPG framework for the design of airfoil.
4.1 Actor and critic neural networks
Figure 7: Actor network architecture
@@ 4.1 Actor and critic neural networks
The actor, as shown in fig. 7, is developed to choose the potential action based on the input state such that the obtained reward is maximal. While the state is the tensor representation of the corresponding airfoil shape, the actor should capture the main feature and distinguish different shapes. Aiming at having a precise control of the airfoil, the geometry is constructed with hundreds of points pair. However, the agent may fail to learn valuable actions based on different states when this tensor representation is conducted within a high-dimensional space. A new state conducted by a specific action only takes a small change of target location. Such a difference is easy to overlook. The smooth action process explained in the previous section can remit such influence but the essential difficulties have not been solved. Trying to conduct the deformation within a lower dimensional tensor space, reduction module should be constructed.
The variational autoencoder meets the requirement of dimensionality reduction and we tested the integration at first. However, it brings extra problems to the construction of the learning process. Firstly, it is not easy to take the balance of the reconstruction error and original loss function. Secondly, deeper neural networks aggravate the risk of the gradient disappearance problem. Thirdly, the sampling process is expensive in this reinforcement learning task. The VAE module can not perform well with limited data and gets trapped into overfitting issues.
@@
Then we tried a different method, using Bézier curves representation for the dimensionality reduction. The Bézier curves take the most important information for specific curves. Besides, the difference between different control points originating from corresponding states is amplified with the least square process. It helps the agent to distinguish different states.
Figure 8: Critic network architecture
@@
The dimension is reduced efficiently with the Bézier curve representation. Then the information should be processed through the neural network to obtain the action that maximizes the cumulative reward from the current state. The output includes action probability for different types of action and corresponding parameters that action possesses. Action that is chosen to be performed originates from the $softmax$ layer while the corresponding parameters are based on the restriction of shape design. For fear of mesh tangling, an activation function needs to be designed that maps the parameters to reasonable intervals. However, the traditional activation functions like $tanh$ and $sigmoid$ suffer from the gradient disappear phenomenon. Then the $softsign$ type activation function is adopted. Techniques like residual connections[47], the SELU activation function[48], and real-time monitoring via TensorBoard[49] were employed to mitigate and monitor potential gradient vanishing problems.
@@
The critic neural network, as illustrated in fig. 8, is designed to evaluate the potential feedback for various actions acted on different states based on the experience accumulated in the replay buffer. The reward mechanism is complicated since the data originate from a CFD solver, not even to mention the shape deformation. To approximate such a process, the interaction between the shape, actions and corresponding parameters should be modeled efficiently. We consider adopting the self-attention and cross-attention mechanism together to investigate the inherent relation. The self-attention mechanism can detect those features which take a significant influence on the target functional. Meanwhile, the cross-attention mechanism is adopted to approximate the shape deformation process. It tries to understand how the action with corresponding parameters influenced the shape deformation. The state, action, and action parameters are mapped to the hidden layer with the same dimension to ensure the cross-attention layer can be modeled.
@@
The loss function is an important component which decides the update direction of parameters in neural networks. For different modules, the loss function is defined according to the features. To choose the action that maximizes the reward based on specific states, the agent should evaluate the potential feedback correctly at first. It indicates that the critical neural network should be updated more frequently than the actor neural network. The long-term reward loss is defined as
@@ 4.2 Exploration and exploitation
Exploration and exploitation are fundamental components of the interaction process. Balancing these two components is vital in reinforcement learning. Excessive exploration is time-consuming. On the other hand, too much exploitation might cause missing out on better strategic options. In this framework, we adopt the $\epsilon$-greedy strategy. This is a popular technique that chooses an action at random with probability $\epsilon$ and selects the action that is predicted to have the highest reward based on its current knowledge with probability $1-\epsilon$. Besides, in our framework, we integrate Ornstein-Uhlenbeck (OU) noise as a complementary tool to the $\epsilon$-greedy strategy, specifically targeting the exploration aspect of reinforcement learning in continuous action spaces. While the $\epsilon$-greedy strategy efficiently balances exploration and exploitation by randomly selecting actions with a probability $\epsilon$, OU noise brings more possibility to the exploration process.
@@
Even though the optimization mechanism can be well implemented with the exploration and exploitation process, the learning effect is not ideal usually. The difficulties originate from the complex mechanisms involved in airfoil design. Firstly, it is not easy to identify the correct reward direction in random exploration. Excessive penalization can lead the neural network parameters into an unfavorable bottleneck, where the agent opts for minimal action to avoid penalties, eventually leading to a suboptimal local optimum. Secondly, airfoil design inherently demands a geometry with a certain level of smoothness. If the shape changes drastically in a single step, even a correct design trend may result in unfavorable reward feedback due to geometric roughness. This also indicates that the shape optimal design needs to be implemented within an elaborate parameter representation environment.
@@
To address the above issues, we introduced several improvement strategies, i). if a certain action results in a penalty, the agent attempts to explore the opposite direction starting from the pre-deformation geometry. Both exploration results and parameters are stored in the replay buffer, and the action leading to a better outcome is chosen for the next step. This mechanism significantly reduces aimless exploration in airfoil design, preventing the model from extensively exploring in unproductive directions. This is also inspired by the idea of a twin network to prevent oscillation during the simulation and it turned out to be very effective for reinforcement learning. Similarly, if the current action brings a perfect reward, the agent will be encouraged to take a similar action with a tiny noise. As the environment is not straightforward to obtain positive meaningful rewards, such a mechanism can help the agent to accumulate useful experience quickly. ii). the single step movement of the sample points is constrained, so that there may not occur irregular geometry. Besides, an early stop mechanism is constructed when the accumulated reward decreases to the minimum threshold. This can be explained as the agent failing for the current design process when the geometry gets trapped into a very terrible shape and may not recover easily. Then the environment should be reset for the next turn. iii). the training process is separated into different periods. A small batch size is adopted for the initial stage. When the replay buffer accumulates enough data, a larger batch size number will be updated. Besides, the data are stored in the replay buffer with a 4-tuple form transition $(state,action,action\_parameters,next\_state)$. Every time the gradient should be calculated from the 4-tuple with a number of batch sizes. The components include three parts, the recent transitions, the transitions with the best reward and randomly chosen transitions from the whole memory. The recent transitions represent the behavior under the current parameters of neural networks while the best transitions can help the agent to repeat the beneficial actions and the random transitions can make the gradient calculation more stable. With such efficient exploration and exploitation mechanisms, we step further to fulfill the reinforcement learning framework.